# Decoding Neural Emotion Patterns through Natural Language Processing Embeddings

### v1

@@ Abstract
Introduction
Understanding the neural correlates of emotional expression in natural language represents a significant challenge in computational neuroscience and affective computing. While traditional neuroimaging studies require expensive equipment and controlled laboratory environments, the increasing availability of digital text data presents new opportunities for emotion-brain mapping. Previous research has primarily focused on either neuroimaging-based emotion localization or computational text analysis independently, with limited integration between these domains. This study proposes a novel computational approach that, while not validated against direct neuroimaging, explores potential relationships between textual emotional content and anatomically-defined brain regions.
Methods
We developed a computational pipeline that maps textual emotion expressions to specific brain regions through a multi-stage process. The framework utilizes OpenAI’s text-embedding-ada-002 model to generate high-dimensional semantic representations of input texts, followed by dimensionality reduction and clustering to identify emotional clusters. These clusters are then mapped to 18 predefined neuro-anatomic brain regions associated with emotional processing. Three distinct experiments were conducted. First, we analyzed conversational data from healthy and depressed subjects using the Distress Analysis Interview Corpus/Wizard-of-Oz (DIAC-WOZ) dataset, comparing emotional brain mapping patterns between these populations. Next, we repeated this process the comprehensive GoEmotions classification dataset. Emotional intensity was quantified using a lexical scoring system that evaluates keyword presence, syntactic patterns, and linguistic modifiers. Finally, we performed a comparison between human-produced text and the response generated by a Large Language Model (LLM) chat bot to evaluate differences in emotional brain activation patterns and determine the extent to which AI-generated language mirrors human emotional expression.
Results
Our proposed approach successfully mapped textual emotions to neuro-anatomically plausible brain regions with high spatial specificity across experiments. Distinct activation patterns emerged between healthy and depressed populations, with depressed subjects showing increased engagement of limbic regions associated with negative affect processing. Extended emotion analysis demonstrated successful differentiation of discrete emotional states. Our final experimental results revealed that while LLM-generated responses exhibited a similar distribution of basic emotions, they lacked the nuanced regional activation observed in human text, particularly in areas associated with empathy and self-referential processing such as the medial prefrontal cortex and posterior cingulate cortex.
Conclusion
This study presents a computational framework for directly mapping natural language emotional content to brain regions without requiring neuroimaging data. The novel integration of semantic embeddings, unsupervised clustering, and neuro-anatomical mapping provides a scalable approach for emotion-brain research that can process large-scale textual datasets. The framework’s ability to distinguish between healthy and depressed populations, as well as differentiate among discrete emotions, demonstrates its potential for both clinical applications and basic emotion research. The methodology offers significant advantages over traditional neuroimaging approaches, including cost-effectiveness, scalability, and the ability to analyze naturalistic language data. Finally, this framework provides a brain-inspired benchmark for evaluating how closely AI-generated language mirrors human emotional expression by comparing their inferred neural activation patterns.
@@ Conclusion
journal: Computers in Biology and Medicine
@@
The integration of semantic embeddings and neuro-anatomical mapping successfully differentiated between healthy and depressed populations through distinct limbic activation patterns.
The framework demonstrated high spatial specificity by accurately mapping twenty-seven discrete emotions to neuro-anatomically plausible brain regions.
@@
In favor of reproducible research and to advance the field, all programming code used in this study is made publicly available.
@@ 1 Introduction
Understanding the neural correlates of emotion has traditionally relied on neuroimaging modalities such as electroencephalography (EEG) and functional magnetic resonance imaging (fMRI) [1, 2, 3]. These techniques have revealed the involvement of regions like the amygdala, insula, anterior cingulate cortex, and prefrontal cortex across different emotional states [4]. Meta-analyses of neuroimaging studies have consistently identified these key regions across diverse emotional paradigms, with the amygdala showing particular importance for threat detection and fear processing [5], while the anterior cingulate cortex and insula demonstrate critical roles in emotional awareness and interoceptive processing [6, 7]. However, traditional neuroimaging approaches face significant limitations including high costs, restricted accessibility, controlled laboratory requirements, and limited ecological validity when studying naturalistic emotional expression [8, 9]. The increasing availability of digital text data presents unprecedented opportunities for emotion-brain mapping that could overcome these existing constraints.
@@
This represents a significant gap in our ability to study emotional processing at scale. The motivation for developing a purely computational emotion–brain mapping framework therefor stems from several converging factors. First, the exponential growth of digital text data offers an unprecedented window into human emotional expression in naturalistic contexts [20]. Second, advances in semantic embedding technologies have created powerful tools for capturing nuanced relationships between language, meaning, and affective content. Third, decades of neuroimaging research have established a robust understanding of the neuro-anatomical basis of emotion processing, with meta-analytic evidence highlighting key neural circuits involved in emotional regulation and expression [18, 21].
The theoretical foundation of our proposed approach rests on the principle that emotional expression in language reflects underlying neural processes. This hypothesis is supported by evidence showing that the human brain encodes language through distributed, continuous representations [22, 23, 24]. Recent studies demonstrated a direct alignment between high-dimensional language model embeddings used by LLMs and neural activation patterns in language-processing regions [23, 24, 25]. Such findings suggest that embedding spaces learned by LLMs may mirror the brain’s own semantic encoding mechanisms, indirectly capturing semantic and emotional features that align with patterns of neural activity.
Furthermore, individuals with certain mental health conditions may exhibit distinct language patterns, including variations in word choice, emotional tone, and syntactic complexity that correlate with well-documented neuro-biological abnormalities [26, 27, 8, 28, 29, 30, 31]. Extending this perspective, characteristic language patterns in clinical populations may manifest as distinct clusters within embedding space, reflecting altered cognitive and affective processing [32, 33, 34, 30]). If these deviations also correspond to measurable shifts in brain activation, embedding-based models could offer novel insights into brain-language relationships and help inform computational mental health diagnostics.
@@
This study therefor investigates whether it is possible to computationally map natural language emotional expressions directly to brain regions without neuroimaging data, using semantic embeddings and clustering techniques to bridge text analysis with neuro-anatomical mapping. Specifically, we aim to:
@@
    Develop a novel computational framework that transforms textual emotional content into neuro-anatomically plausible brain region activations using state-of-the-art natural language processing techniques.
@@
    Provide an objective, brain-based measure of human-likeness in language by comparing AI-generated and human-authored texts through their inferred emotional brain activation patterns.
@@ 2.1 Datasets, Preprocessing and Text Preparation
2.2 Text Embedding Generation
During preprocessing (Figure 1, step 1), text was segmented into chunks of approximately 300 characters using periods as sentence boundaries. Next, text embeddings were generated using OpenAI’s [40] text-embedding-ada-002 model (Figure 1, step 2), which produces 1536-dimensional vector representations of input text. This model was chosen for its strong performance in capturing both semantic relationships and emotional nuances in natural language [28, 29, 30, 31]. Although training a custom embeddings model for our experimentation is technically feasible, it could introduce unwanted bias into our experiments, while the OpenAI text-embedding-ada-002 embeddings are commonly used in many public and commercial LLMs.
Figure 1: Five-step text computational pipeline to convert natural language text to embeddings, reduce dimensionality, cluster to emotional groups and map to final brain regions.
2.3 Dimensionality Reduction and Spatial Mapping
2.4 Emotional Intensity Estimation
  • 1.
  • 2.
  • 3.
  • 4.
  • 1.
  • 2.
  • 3.
    Exclamation marks added 0.25 per occurrence (maximum 4)
  • 4.
    Question marks contributed 0.15 per occurrence (maximum 3)
  • 5.
    All-caps text (more than 3 characters) added 0.5 intensity
Base intensity was set at 0.1 for all texts to account for implicit emotional content, with final intensity scores capped at 2.0 to prevent extreme outliers from skewing subsequent analyses.
2.5 Emotion Region Clustering
K-means clustering was applied to the 3D PCA-transformed embeddings to identify distinct emotional patterns within the data (Figure 1, step 4). The number of clusters was set to match the number of predefined anatomical brain regions (11 bilateral + 7 midline = 18 regions), establishing a direct correspondence between emotional content clusters and neuro-anatomical structures [5, 1, 47, 48, 49]. Of the 18 anatomically defined brain regions selected, 14 have been consistently implicated in emotion processing [3, 50, 2, 1].
2.6 Cluster-to-Region Assignment
The assignment of emotional content clusters to specific brain regions (Figure 1, step 5) employed a two stage distance minimization approach that preserved the one-to-one mapping between clusters and anatomical regions.
First, each cluster center’s coordinates in the 3D PCA space were compared against the predefined Montreal Neurological Institute (MNI) coordinate positions of the 18 emotion processing brain regions using Euclidean distance. These regions were selected to span key nodes of the emotional circuitry and included limbic structures, prefrontal regions, subcortical structures, temporal regions, and brain stem nuclei, along with anterior and posterior cingulate cortex and medial prefrontal cortex.
This sequential assignment continued until all clusters were mapped to unique brain regions, ensuring that the spatial distribution of emotional content in the 3D space corresponded meaningfully to the anatomical organization of emotion-processing brain networks. Once cluster-to-region assignments were established, individual text samples inherited their brain region labels based on their cluster membership, creating the final mapping from textual content to neuro-anatomical locations.
2.7 Statistical Analysis
The computational pipeline incorporated statistical practices including random seed setting to ensure reproducibility and management of edge cases such as insufficient sample sizes. Region-specific analysis was conducted by aggregating texts assigned to each brain region and calculating mean emotional intensities, providing quantitative measures of regional emotional activation patterns. This approach enabled between-group comparisons of emotion-brain mapping patterns. Figure 2 provides a visual hierarchy detailing the emotion to brain region mapping approach as detailed in steps 4 and 5 of Figure 1.
The programming pseudo-code for the three key algorithms utilized and described in Section 2 are detailed below:
@@ 2.7 Statistical Analysis
Algorithm 3 Emotion Region Clustering and Brain Assignment
@@
3:  $regions\leftarrow$ {25 brain regions with MNI coordinates}
@@
7:function PerformClustering(embeddings_3dembeddings\_3d, n_regions=25n\_regions=25)
@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Emotion mapping results from the first experiment revealed notable differences in neural activity patterns between clinical interview transcripts of healthy individuals and those with depression (Figure 3). The analysis shows distinct activation profiles across different brain regions when comparing healthy to depressed subjects. For the healthy individuals, robust emotional activations were observed across multiple brain regions, with particularly strong responses in several key areas. The insula showed the highest activation levels (reaching 7 units), followed by the isthmus region (also 7 units), and the pericalcarine cortex (7 units). Other regions showing substantial activation included the amygdala (6 units), anterior cingulate (3 units), and hippocampus (6 units) [6]. This widespread activation pattern suggests healthy emotional processing involves coordinated activity across multiple brain networks, particularly within limbic-cortical circuits.
In contrast, the depressed group demonstrated a markedly different activation profile characterized by consistently lower activation levels across nearly all brain regions examined. Most regions in the depressed group showed activation levels between 1-3 units, representing substantial reductions compared to healthy controls. The most pronounced differences were observed in key emotional processing regions including the insula which dropped from 7 to approximately 2 units, the isthmus (reduced from 7 to 2 units), and the pericalcarine cortex (reduced from 7 to 1 unit [51]).
Figure 3: Comparison of activation counts per brain region for healthy versus depressed subjects.
The broader analysis by brain system categories (Figure 4) revealed systematic differences in activation patterns. Cortical regions showed the most substantial difference, with healthy subjects displaying approximately 40 total activations compared to about 13 in depressed subjects (a 67% reduction). Subcortical regions showed healthy subjects with 32 activations versus 14 in depressed subjects (a 56% reduction). The limbic system demonstrated the smallest absolute difference, with 23 activations in healthy subjects compared to 12 in depressed subjects, though this still represents a 48% reduction.
Figure 4: Brain region analysis for healthy versus depressed subjects.
Figure 5 presents a 3D cortical surface rendering in MNI space, displaying spatially localized differences in emotional processing between healthy and depressed subject groups (Table 2). The visualization employs a heat map approach where red regions indicate areas of greatest divergence in emotional activation patterns between the two populations, with intensity values ranging from 0.00 to 1.00 as shown in the color scale.
The rendering reveals distinct anatomical clusters of emotional processing differences, with the most pronounced activation disparities concentrated in several key regions. Notable high-intensity areas (approaching the maximum 1.00 value) are observed in bilateral insula regions, particularly prominent in the left hemisphere, along with posterior cingulate cortex involvement and brain stem areas corresponding to the raphe nuclei complex. Additional moderate-intensity differences are visible in frontal and temporal cortical regions. The left insula is associated with heightened perception of internal states, while raphe nuclei hyperactivity may indicate dysregulated serotonergic firing, leading to ineffective emotional modulation [55, 56, 57].
Figure 5: 3D rendering of the caudal (left) and parietal (right) regions with emotion intensity differences (Table 2) healthy and depressed subjects shown in red.

### v2

# Decoding Neural Emotion Patterns through Large Language Model Embeddings

@@ Abstract
Linking emotional expression in natural language to brain function remains a central challenge in neuroscience and affective computing. While neuroimaging offers insight into emotion localization, it is costly and limited to controlled settings. In this study, we introduce a computational framework that maps textual emotional content to neuroanatomical brain regions, enabling scalable, imaging-free emotion–brain research. Our proposed method combines semantic embeddings, dimensionality reduction, and clustering to identify emotional states, which are mapped onto brain regions associated with affective processing. We applied this framework across three experiments: i) comparing conversational data from healthy and depressed individuals, ii) analyzing a large-scale emotion dataset, and iii) contrasting human language with outputs from a large language model. Emotional intensity was quantified using a lexical scoring system sensitive to keywords, syntax, and modifiers. This approach produced neuro-anatomically plausible mappings with high spatial specificity. Depressed subjects exhibited reduced engagement of the left insula and raphe nuclei, regions linked to interoceptive awareness and serotonergic function. The framework reliably differentiated discrete emotions and revealed that LLM-generated texts, while capturing basic affective patterns, lacked the nuanced activation seen in human language, particularly in empathy and self-referential regions such as the medial prefrontal and posterior cingulate cortices. This work establishes a scalable, cost-effective alternative to neuroimaging, advancing both clinical and computational models of emotion, while providing a neuro-inspired benchmark for assessing how closely AI-generated language mirrors human emotional expression.
@@
journal: Neurocomputing
@@
The integration of semantic embeddings and neuro-anatomical mapping successfully differentiated between healthy and depressed populations through distinct activation patterns.
The framework demonstrated high spatial specificity by accurately mapping discrete emotions to neuro-anatomically plausible brain regions.
@@
In favor of reproducible research and to advance the field, all programming code/data used in this study is made publicly available.
@@ 1 Introduction
Understanding the neural correlates of emotion has traditionally relied on neuroimaging modalities such as electroencephalography (EEG) and functional magnetic resonance imaging (fMRI) [1, 2, 3]. These techniques have revealed the involvement of regions like the amygdala, insula, anterior cingulate cortex, and prefrontal cortex across different emotional states [4]. Meta-analyses of neuroimaging studies have consistently identified these key regions across diverse emotional paradigms, with the amygdala showing particular importance for threat detection and fear processing [5], while the anterior cingulate cortex and insula demonstrate critical roles in emotional awareness and interoceptive processing [6, 7]. However, traditional neuroimaging approaches face significant limitations, including high costs, restricted accessibility, controlled laboratory requirements, and limited ecological validity when studying naturalistic emotional expression [8, 9]. The increasing availability of digital text data presents unprecedented opportunities for emotion-brain mapping that could overcome these existing constraints.
@@
This represents a significant gap in our ability to study emotional processing at scale. The motivation for developing a purely computational emotion–brain mapping framework, therefore, stems from several converging factors. First, the exponential growth of digital text data offers an unprecedented window into human emotional expression in naturalistic contexts [20]. Second, advances in semantic embedding technologies have created powerful tools for capturing nuanced relationships between language, meaning, and affective content. Third, decades of neuroimaging research have established a robust understanding of the neuro-anatomical basis of emotion processing, with meta-analytic evidence highlighting key neural circuits involved in emotional regulation and expression [18, 21].
The theoretical foundation of our proposed approach rests on the principle that emotional expression in language reflects underlying neural processes. This hypothesis is supported by evidence showing that the human brain encodes language through distributed, continuous representations [22, 23, 24]. Studies have demonstrated a direct alignment between high-dimensional language model embeddings and neural activation patterns in language-processing regions [23, 24, 25].
Recent work by Goldstein et al. [26] further established that simple linear encoding models can reliably map language model embeddings onto brain activity during natural conversation, achieving robust correlations with recorded signals. Their findings showed that deep embeddings consistently outperform traditional symbolic linguistic features (e.g., phonemes, syntactic categories) in predicting neural responses across cortical regions, suggesting that high-dimensional continuous representations more faithfully reflect the brain’s internal language mechanisms. Importantly, they also revealed individual and regional differences in neural–embedding alignment, underscoring the capacity of computational models to capture variability in neural organization that extends beyond region-specific specialization. Such evidence not only validates embedding-based approaches for brain-language mapping but also provides a methodological and theoretical foundation for extending this framework to emotion, where capturing both shared and individual patterns of neural processing is particularly critical.
Furthermore, individuals with certain mental health conditions may exhibit distinct language patterns, including variations in word choice, emotional tone, and syntactic complexity that correlate with well-documented neuro-biological abnormalities [27, 28, 8, 29, 30, 31, 32]. Extending this perspective, characteristic language patterns in clinical populations may manifest as distinct clusters within embedding space, reflecting altered cognitive and affective processing [33, 34, 35, 31]. If these deviations also correspond to measurable shifts in brain activation, embedding-based models could offer novel insights into brain-language relationships and help inform computational mental health diagnostics.
@@
This study, therefore, investigates whether it is possible to computationally map natural language emotional expressions directly to brain regions without neuroimaging data, using semantic embeddings and clustering techniques to bridge text analysis with neuro-anatomical mapping. Specifically, we aim to:
@@
    Develop a novel computational framework that transforms textual emotional content into neuro-anatomically plausible brain region activations using state-of-the-art natural language processing embedding techniques.
@@
    Provide an objective, brain-based measure of human-likeness in language by comparing LLM-generated and human-authored texts through their inferred emotional brain activation patterns.
@@ 2.1 Datasets
2.2 Text Preprocessing and Embedding Generation
During preprocessing (Figure 1, step 1), text was segmented into chunks of approximately 300 characters using periods as sentence boundaries. Next, text embeddings were generated using OpenAI’s [41] text-embedding-ada-002 model (Figure 1, step 2), which produces 1536-dimensional vector representations of input text. This model was chosen for its strong performance in capturing both semantic relationships and emotional nuances in natural language [29, 30, 31, 32]. Although training a custom embeddings model for our experimentation is technically feasible, it could introduce unwanted bias into our experiments, while the OpenAI text-embedding-ada-002 embeddings are commonly used in many public and commercial LLMs. The programming pseudo-code for steps 1 and 2 of Fig. 1 is shown in Algorithm 1.
Figure 1: Five-step computational pipeline to convert natural language text to embeddings, reduce dimensionality, cluster to emotional groups and map to brain regions.
@@ 2.2 Text Preprocessing and Embedding Generation
2.3 Dimensionality Reduction and Spatial Mapping
2.4 Emotional Intensity Estimation
  • 1.
  • 2.
  • 3.
  • 4.
  • 1.
  • 2.
  • 3.
    Exclamation marks added 0.25 per occurrence (maximum 4)
  • 4.
    Question marks contributed 0.15 per occurrence (maximum 3)
  • 5.
    All-caps text (more than 3 characters) added 0.5 intensity
Base intensity was set at 0.1 for all texts to account for implicit emotional content, with final intensity scores capped at 2.0 to prevent extreme outliers from skewing subsequent analyses. The programming pseudo-code for steps 3A and 3B of Fig. 1 is shown in Algorithm 2.
@@ 2.4 Emotional Intensity Estimation
2.5 Emotion Region Clustering
K-means clustering was applied to the 3D PCA-transformed embeddings to identify distinct emotional patterns within the data (Figure 1, step 4). The number of clusters was set to match 29 predefined anatomical brain regions (Table 2), establishing a direct correspondence between emotional content clusters and neuro-anatomical structures [5, 1, 48, 49, 50]. Of the 29 anatomically defined brain regions selected, 14 have been consistently implicated in emotion processing [3, 51, 2, 1].
Table 2: Predefined anatomical brain regions (29) used for k-means clustering.
Anatomical Category Brain Regions
Limbic System Amygdala Left, Amygdala Right, Hippocampus Left, Hippocampus Right,
Anterior Cingulate Left, Anterior Cingulate Right
Cortical Regions Prefrontal Cortex Left, Prefrontal Cortex Right, Medial Prefrontal Cortex,
Orbitofrontal Left, Orbitofrontal Right, Insula Left, Insula Right,
Temporal Pole Left, Temporal Pole Right, Superior Temporal Left,
Superior Temporal Right, Posterior Cingulate
Subcortical Regions Caudate Left, Caudate Right, Putamen Left, Putamen Right,
Nucleus Accumbens Left, Nucleus Accumbens Right, Hypothalamus,
Periaqueductal Gray, Ventral Tegmental Area, Raphe Nuclei,
Locus Coeruleus
2.6 Cluster-to-Region Neuroanatomical Mapping
The assignment of emotional content clusters to specific brain regions (Figure 1, step 5) employed a two-stage distance minimization approach that preserved the one-to-one mapping between clusters and anatomical brain regions.
First, each cluster center’s coordinates in the 3D PCA space were compared against the predefined Montreal Neurological Institute (MNI) [52] coordinate positions of the 29 emotion processing brain regions using Euclidean distance. These regions were selected to span key nodes of the emotional circuitry and included limbic structures, prefrontal regions, subcortical structures, temporal regions, and brain stem nuclei, along with anterior and posterior cingulate cortex and medial prefrontal cortex.
This sequential assignment continued until all clusters were mapped to unique brain regions, ensuring that the spatial distribution of emotional content in the 3D space corresponded meaningfully to the anatomical organization of emotion-processing brain networks. Once cluster-to-region assignments were established, individual text samples inherited their brain region labels based on their cluster membership, creating the final mapping from textual content to neuro-anatomical locations. Figure 2 provides a visual explanation of this cluster to brain region mapping process.
Figure 2: Text embedding clusters mapped to brain regions via PCA dimension reduction based on neuro-scientifically plausible regions.
The programming pseudo-code for steps 4 and 5 of Fig. 1 is shown in Algorithm 3. Figure 3 provides a visual hierarchy detailing the emotion to brain region mapping approach as detailed in steps 4 and 5 of Figure 1.
Algorithm 3 Emotion Region Clustering and Brain Region Assignment
@@ 2.6 Cluster-to-Region Neuroanatomical Mapping
3:  $regions\leftarrow$ {29 brain regions with MNI coordinates}
@@
7:function PerformClustering(embeddings_3dembeddings\_3d, n_regions=29n\_regions=29)
@@
2.7 Statistical Analysis
The proposed computational pipeline incorporated statistical practices, including random seed setting to ensure reproducibility and management of edge cases such as insufficient sample sizes. Region-specific analysis was conducted by aggregating texts assigned to each brain region and calculating mean emotional intensities, providing quantitative measures of regional emotional activation patterns. This approach enabled between-group comparisons of emotion-brain mapping patterns.
2.8 Multi-Trial Validation and Clustering Pattern Analysis
To validate the robustness of our findings and address potentially confounding statistical interpretations due to the differences in subject counts (97 healthy vs 37 depressed), we conducted a comprehensive multi-trial analysis consisting of 15 independent trials across three class balancing strategies: i) undersampling through randomly downsampled healthy subjects to match the minority class (n=37 per group), ii) oversampling depressed subjects with replacement to match the majority class (n=97 per group), and iii) a hybrid strategy which resampled both groups to an intermediate size (n=67 per group). Prior to the main analysis, ablation studies using a subset of healthy participants (n=50) evaluated dimensionality reduction methods (PCA vs. t-SNE), with PCA achieving optimal silhouette scores across all tested configurations.
Each of the 15 trials utilized a unique random seed to ensure reproducible yet independent sampling, with the identical five-step computational pipeline (Figure 1) applied to all balanced datasets. Bootstrap confidence intervals were calculated using 50 resamples per trial, and clustering quality was assessed using silhouette scores as the primary outcome measure. Mann-Whitney U tests evaluated statistical significance of intensity differences between groups, while Cohen’s d estimated effect sizes using pooled standard deviations derived from confidence intervals.
Quality control measures included k-means convergence verification, embedding consistency validation, and spatial mapping checks to ensure one-to-one correspondence between emotion clusters and brain regions. Results were aggregated across trials to compute mean p-values, percentage of significant trials, effect size distributions, and clustering quality ratios (depressed silhouette scores vs. healthy silhouette scores).
@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Emotion mapping results from the first experiment performed on the DIAC-WOZ dataset [37] that comprises annotated interview transcripts from individuals diagnosed with depression and healthy controls revealed notable differences in neural activations between healthy individuals and those with depression (Figure 4). The analysis shows distinct activation profiles across different brain regions (as categorized in Table 2) when comparing healthy to depressed subjects. For the healthy individuals, activations were observed across multiple brain regions, with particularly strong responses in two key areas. The insula (located in the cortical region) [6, 53, 54] and raphe nuclei (located in subcortical region) [55, 56, 57] showed the highest activation levels that were statistically significant.
Figure 4: Comparison of activation count differences per brain region for healthy versus depressed subjects.